OpenStudy (glorylove):

So I'm having such a hart time with simple interest! I understand the formula is I = prt, but I'm still so confused on this problem. I don't need the answer given to me, just the means on how to find it please! Zach borrowed $4,000 from his grandfather to buy a used car so he could travel to and from work. After working for 3 years, he has saved $5,440 to pay back to his grandfather. At what interest rate did Zach borrow the money? Could someone please help me with finding the interest rate?

jimthompson5910 (jim_thompson5910):

He borrowed $4,000, so P = 4000 The amount he pays back is assumed to be $5,440 So the interest he has paid is 5440 - 4000 = 1440 dollars The interest rate is unknown. We'll just leave it as 'r' The time is 3 years, so t = 3

jimthompson5910 (jim_thompson5910):

I = Prt 1440 = 4000*r*3 ... plug in the given values (see above) 1440 = 4000*3*r 1440 = 12000r 1440/12000 = r 0.12 = r r = 0.12 In percentage form, the interest rate is 12% (move the decimal over 2 spots to the right)

OpenStudy (glorylove):

Oh my goodness, you're a miracle solver. I think I was trying to overcomplicated it in my mind. Thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (glorylove):

*over-complicate.. man. Whoops. But yah get it.